# Towards Exploring Continual Fine-Tuning for Enhancing Language Ability in Large Language Models

### v1

@@ Abstract
A common challenge towards the adaptability of Large Language Models (LLMs) is their ability to learn new languages over time without hampering the model’s performance on languages in which the model is already proficient (usually English). Continual fine-tuning (CFT) is the process of sequentially fine-tuning an LLM to enable the model to adapt to downstream tasks with varying data distributions and time shifts. This paper focuses on the language adaptability of LLMs through CFT. We study a two-phase CFT process in which an English-only end-to-end fine-tuned LLM from Phase 1 (predominantly Task Ability) is sequentially fine-tuned on a multilingual dataset – comprising task data in new languages – in Phase 2 (predominantly Language Ability). We observe that the “similarity” of Phase 2 tasks with Phase 1 determines the LLM’s adaptability. For similar phase-wise datasets, the LLM after Phase 2 does not show deterioration in task ability. In contrast, when the phase-wise datasets are not similar, the LLM’s task ability deteriorates. We test our hypothesis on the open-source Mistral-7B and LLaMA-3-8B models with multiple phase-wise dataset pairs. To overcome the deterioration, we propose two methods, based on layer freezing and generative replay. We show that our methods assist the LLM in improving its language ability while preserving its task ability.
@@ 1 Introduction
Figure 1: Comparing hidden activations for Mistral-7B and LLaMA-3-8B during our two-phase continual fine-tuning process. We prompt each model with examples from MTBench (zheng2024judging), and visualize the similarity between the mean hidden activations, for each model layer. For datasets that encode “similar” tasks (Alpaca & MultiAlpaca), model’s task ability does not decline (e.g., 3% gain for IFEval). For non-similar datasets (Instruct & MultiAlpaca), the task ability declines (e.g., 8% decline for IFEval). Here, Phase 2 model representations do not align with Phase 1’s; thus, suggesting greater model weight interference and a decline in task ability.
With ever-increasing adoption of LLMs in real world applications and expanding multilingual user bases of these applications, it is important to cater these models to wide enough multilingual audiences. Model training is compute hungry, and there is an abundance of both labelled and unlabelled data in English as compared to other languages (shaham2024multilingual). As such, it is imperative to find efficient ways to use pre-trained or fine-tuned models to improve performance on other languages. In this paper, we refer to a model’s ability in non-English languages as predominantly its language ability (LA), which can be achieved without relying on large amounts of data in those languages. Instead, we can exploit the predominantly task ability (TA) learned from English data.
To this end, researchers use techniques like continual pre-training, continual fine-tuning or language adaption to adapt models to a newer set of languages to enhance their language abilities (refer to §LABEL:sec::rw). While these techniques are effective, they are highly task-specific. Furthermore, existing techniques for multilingual LLMs rely on parallel data, old fine-tuning data, or old and new set of parameters. Parameter efficient techniques like LoRA (hu2022lora) are also widely used to efficiently fine-tune LLMs on multilingual data. However, such techniques show both: catastrophic forgetting on English and incapability to exploit the task ability that the model receives from the English fine-tuning data (aggarwal2024maple).
In such a setting, we want to enhance the model’s language ability (other than English) while preserving the task ability achieved via (firstly) English fine-tuning. This setting results in the challenge of catastrophic forgetting, i.e., the model’s task ability on English may decline while fine-tuning on multilingual data (mukhoti2023fine). Furthermore, a trivial solution that fine-tunes on the mixture of multilingual and English-only data may be sub-par (e.g., due to language relatedness (dhamecha2021role)). Hence, it is challenging to improve an LLM’s language ability while preserving its performance on English.
Our Approach.
We use a two-phase continual fine-tuning (CFT) technique for language adaption. We study the effects of various English and multilingual instruction tuning datasets when the models are fine-tuned in two phases: where Phase 1 is fine-tuning the model in English to improve its task ability and then fine-tuning it on a proportionally-sized multilingual dataset in Phase 2. In Phase 1, we use Alpaca (alpaca) and OpenOrca (OpenOrca), and in Phase 2 we use MultiAlpaca (multialpaca) and mOpenOrca (§3.1).
@@
We perform this study on two open-source models, namely LLaMA-3-8B and Mistral-7B. We also use fine-tuned versions of them, namely LLaMA-3-8B-Instruct and Mistral-7B-Instruct, as off-the-shelf Phase 1 fine-tuned models. We quantify a model’s task ability based on its performance on four English datasets: (i) two for instruction following (i.e., IFEval (zhou2023instruction) and Alpaca Eval (li2023alpacaeval)) and (ii) two for reasoning tasks (i.e., MMLU (hendryckstest2021) and HellaSwag (zellers2019hellaswag)). Likewise, we quantify a model’s language ability based on its performance on (i) two question answering tasks (i.e., MLQA (lewis2019mlqa) and XQuAD (Artetxe:etal:2019)) and (ii) XLSUM (hasan-etal-2021-xl), a summarization task.
Our Contributions.
@@
First, we observe that when phase-wise English and multilingual datasets encode different tasks, we see a decline in the Phase 2 model’s performance on English. On the other hand, when Phase 1 and Phase 2 datasets encode similar tasks, the Phase 2 model’s performance on English improves (refer to Figure 1). Second, to quantify the similarity of these phase-wise datasets, we introduce two metrics based on language-agnostic embeddings and model representations. We show that our quantification correlates with the decline in task ability (§3.3). Third, we study the efficacy of two tailored variants of existing CFT strategies to mitigate the decline in task ability after Phase 2 fine-tuning, while also boosting the language ability. The first strategy we study is generative replay, i.e., using instructions from a similar English counterpart of the Phase 2 dataset to generate replay data using the Phase 1 model. The second strategy uses heuristic-based layer freezing. Here, we use the weight difference between the Base and Phase 1 models to pick specific layers for freezing during Phase 2 fine-tuning. We study the gains in task and language ability of these strategies compared to specific baselines (§4).
2 Enhancing Language Ability through Continual Fine-tuning
A common recipe to training LLMs to learn new languages is to use a training paradigm that focuses on task and language adaption (chen2024improvinglanguageplasticitypretraining). Concretely, we define task adaption as the model’s ability to comprehend the input text and then provide a suitable output. We refer to language adaption as the model’s ability to perform those tasks in languages other than English.
In Task Adaption, the LLM is trained to follow instructions, usually using labeled English data. Language Adaption focuses on training the LLM to understand text from newer languages. Task adaptation leverages cross-lingual transfer, facilitating language adaptation to a certain degree. However, this process can result in a decline in the LLM’s task adaptation performance due to the risk of catastrophic forgetting. Despite this challenge, task adaptation often yields greater benefits compared to relying solely on cross-lingual transfer for language adaptation.
Continual Fine-tuning for Language Adaption.
To improve the language adaption of LLMs, we re-imagine the above recipe as a two-phase CFT process. We have:
  • Phase 1. We fine-tune a base LLM end-to-end on an English instruction dataset. Phase 1 aims to predominantly teach the LLM instruction following ability, which we refer to as task ability.
@@ 2 Enhancing Language Ability through Continual Fine-tuning
  • Phase 2. Here, we use the fine-tuned LLM from Phase 1 and further end-to-end fine-tune it on a Multilingual instruction dataset. Unlike chen2024improvinglanguageplasticitypretraining, in our setting, the data in Phase 2 is labeled. However, compared to Phase 1, Phase 2’s dataset is geared towards enhancing the LLM’s language ability, and comprises multiple languages with fewer data points per language.
This paper relates English fine-tuning with task ability enhancement as English fine-tuning predominantly helps in the task ability of LLMs. Whereas multilingual fine-tuning predominantly helps with an LLM’s language ability.
CFT for Language Adaption: Challenges.
@@
The primary challenge in our two-phase fine-tuning process is that the LLM’s language ability must not come at the cost of its task ability. We impose two additional constraints based on real-world scenarios. First, in Phase 2, we cannot re-use Phase 1’s dataset. Often instruction fine-tuned LLMs are available without their corresponding datasets (e.g., Mistral-7B-Instruct (mistral)). Second, in Phase 2, we cannot use the weights of the Phase 1 model during training, as saving both old and new set of parameters on the GPU for training would be computationally expensive.
In a nutshell, we focus on CFT for language adaption for an LLM while preserving the model’s task ability.
@@
Model Phase 1 (P1) Phase 2 (P2) IFEval ()($\uparrow$)( ↑ ) Alpaca Eval ()($\uparrow$)( ↑ ) MMLU ()($\uparrow$)( ↑ ) HellaSwag ()($\uparrow$)( ↑ ) Average Dataset Dataset P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 Mistral-7B Alpaca MultiAlpaca 0.364 0.395 0.12 0.16 0.552 0.573 0.581 0.616 0.404 0.436 Instruct 0.550 0.462 0.35 0.15 0.575 0.533 0.641 0.416 0.529 0.390 LLaMA-3-8B Alpaca 0.277 0.326 0.10 0.11 0.231 0.242 0.556 0.567 0.291 0.311 Instruct 0.735 0.182 0.14 0.10 0.340 0.239 0.533 0.278 0.437 0.2
@@
Table 1: Task Ability results for two-phase Continual Fine-tuning (CFT). When the phase-wise datasets are similar (Definition 1 and Definition 2), task ability post Phase 2 (P2) fine-tuning consistently improves (denoted with green). When the phase-wise datasets are not similar, we see a significant decline in task ability post Phase 2 (P2) fine-tuning (denote with red).
3 Evaluating Task & Language Ability for Multilingual CFT
@@ 3.1 Experiment Setup & Evaluation Tasks
We continually fine-tune open-source Mistral-7B (mistral) and LLaMA-3-8B (llama) LLMs for language adaption.
@@
Fine-tuning Datasets. For our phase-wise datasets, we use the open-source Alpaca (alpaca), MultiAlpaca (multialpaca), and OpenOrca (OpenOrca) datasets. Alpaca is a self-instruct English-only dataset. MultiAlpaca is a multilingual dataset created by translating Alpaca’s seed tasks to 11 languages and using GPT-3.5-Turbo for response collection. The languages are in equal proportions and are “French”, “Arabic”, “German”, “Spanish”, “Indonesian”, “Japanese”, “Korean”, “Portuguese”, “Russian”, “Thai”, and “Vietnamese”. The appendix (§A.2) describes OpenOrca and mOpenOrca.
Fine-tuning Technique. We perform full fine-tuning with bf16 precision to study the effects of full fine-tuning with multilingual data in Phase 2 and its effect on task ability. We also wish to exploit the benefits gained via complete fine-tuning of these models, which may not be possible with parameter efficient fine-tuning (aggarwal2024maple; panda2024lotteryticketadaptationmitigating). However, in §4, we propose a heuristic-based layer freezing strategy to mitigate forgetting of task ability in which we freeze some layers and fine-tune the rest. For our experiments, we use Axolotl222github.com/axolotl-ai-cloud/axolotl/, an open-source framework to fine-tune LLMs. We conducted our experiments on NVIDIA A100 GPUs with 80 GB RAM.
@@
Evaluation Tasks. To quantify an LLM’s task ability, we evaluate Phase 1 and Phase 2 models on two instruction-following tasks (i) IFEval (zhou2023instruction) and (ii) Alpaca Eval (li2023alpacaeval), (iii) MMLU (hendryckstest2021) for problem-solving and (iv) HellaSwag (zellers2019hellaswag) for commonsense reasoning ability. To quantify an LLM’s language ability, we evaluate our fine-tuned models on three benchmark datasets comprising two multilingual generative tasks: question answering (MLQA (lewis2019mlqa) & XQuAD (Artetxe:etal:2019)) and summarization (XLSUM (hasan-etal-2021-xl)). Further details on these tasks are available in the Appendix (§A.3).
@@
To evaluate our models on TA and LA, we use LM-Evaluation-Harness333github.com/EleutherAI/lm-evaluation-harness, which is a unified framework for zero/few-shot evaluations of LLMs. For both task and language ability, we use zero-shot evaluation. For additional details on the training setup, code, and evaluation tasks, we refer the reader to the Appendix (§A).
3.2 Task and Language Ability Results

### v3

# Exploring Two-Phase Continual Instruction Fine-tuning for Multilingual Adaptation in Large Language Models

@@ Abstract
A key challenge for Large Language Models (LLMs) is improving their Multilingual instruction-following ability over time without deteriorating their ability in languages they already excel at, typically English. In this paper, we study a two-phase Continual Fine-tuning (CFT) setup toward improving a model’s Multilingual adaptability. Concretely, we consider a two-phase CFT process in which an English-only end-to-end instruction fine-tuned LLM (Phase 1) is sequentially fine-tuned on a multilingual instruction dataset (Phase 2). Across Mistral-7B and LLaMA-3-8B and multiple dataset pairs, we show that instructional similarity between phases is critical: aligned datasets preserve or improve English while boosting multilingual ability, whereas misaligned datasets cause English degradation. We show that this degradation arises from representation shift during CFT, and that targeted mitigation strategies, including generative replay and heuristic-based layer freezing, reduce this shift and improve multilingual adaptation.
Exploring Two-Phase Continual Instruction Fine-tuning for Multilingual Adaptation in Large Language Models
Divyanshu Aggarwalthanks: Joint First Author1, Sankarshan Damlefootnotemark: 1, Navin Goyal2, Satya Lokam1, Sunayana Sitaram2 1Microsoft  2Microsoft Research India {divyanshu.aggarwal,t-sandamle,navingo,satya.lokam,sunayana.sitaram}@microsoft.com
@@ 1 Introduction
Figure 1: Comparing t-SNEs (van der Maaten and Hinton, 2008) of the hidden activations for Mistral-7B and LLaMA-3-8B during our two-phase Continual Fine-tuning (CFT) process. We prompt each model with examples from MTBench Zheng et al. (2024), and visualize the similarity between the mean hidden activations, for each model layer. For datasets that encode "similar" instructions (Alpaca & MultiAlpaca), English ability does not decline (e.g., 3% gain for IFEval). For non-similar datasets (Instruct & MultiAlpaca), English ability declines (e.g., 8% decline for IFEval). Here, Phase 2 model representations do not align with Phase 1’s; thus, suggesting greater model weight interference and a decline in English ability.
The widespread adoption of Large Language Models (LLMs) has led to a growing multilingual user base Shiyas (2023). However, ensuring strong performance across languages remains a fundamental challenge, with models consistently performing worse on low-resource languages spoken by millions of speakers worldwide (Ahuja et al., 2023, 2024a). A key limitation is that both labeled and unlabeled training data are predominantly available in English and a few high-resource languages, while resources for other languages, especially low-resource ones, are scarce Shaham et al. (2024).
Training large models from scratch is computationally expensive, making fine-tuning pre-trained LLMs the preferred approach for improving multilingual capabilities Lankford et al. (2023); Nguyen et al. (2023). A common fine-tuning strategy is to train LLMs on an instruction-following dataset that contains a mixture of languages. However, these datasets are often heavily skewed toward English and other high-resource languages, leading to a performance imbalance: models perform strongly in English but struggle with low-resource languages Dhamecha et al. (2021); Li et al. (2024a, b). Further, prior works show that fine-tuning on a dataset that only contains non-English languages can hurt the model’s performance on English due to catastrophic forgetting Wu et al. (2025); Behrouz et al. (2025), which is not desirable for most real-world scenarios due to the volume of English queries Ta (2023). Ideally, we want the same model to be proficient in both English and other languages to avoid the costs of maintaining multiple models. We refer to an LLM’s proficiency in English as its English Ability (EA), and its effectiveness across other languages its Multilingual Ability (MA). In this work, we aim to improve an LLM’s MA while maintaining or improving its EA.
1.1 Our Approach
To bridge the gap between EA and MA, we introduce a two-phase Continual Fine-tuning (CFT) setup. We fine-tune a pre-trained LLM on an English instruction dataset in Phase 1 and then fine-tune it on a similarly-sized Multilingual dataset in Phase 2. In Phase 1, we use Alpaca Taori et al. (2023) and OpenOrca Lian et al. (2023), and in Phase 2 we use MultiAlpaca Wei et al. (2023) and mOpenOrca (§4.1). Alpaca and OpenOrca provide high-quality English instruction data, while MultiAlpaca and mOpenOrca are their multilingual counterparts, ensuring consistency in instruction style across phases. To compare the efficacy of our two-phase CFT setup, we compare it with a straightforward single-phase setup where the LLM is fine-tuned on the mixture of both the instruction tuning datasets.
@@ 1.1 Our Approach
We focus on two open-source models, LLaMA-3-8B and Mistral-7B as base models for our experiments. We also use fine-tuned versions of them, LLaMA-3-8B-Instruct and Mistral-7B-Instruct, as off-the-shelf Phase 1 English fine-tuned models [^1] LLaMA-3-8B’s pre-training data was 5% multilingual, but LLaMA-3-8B-Instruct is primarily non-multilingual (Dubey et al., 2024).. We quantify a model’s English Ability (EA) based on its performance on four English datasets: (i) Two datasets that measure instruction following capabilities (i.e., IFEval Zhou et al. (2023) and Alpaca Eval Li et al. (2023)) and (ii) two that measure reasoning abilities (i.e., MMLU Hendrycks et al. (2021) and HellaSwag Zellers et al. (2019)). Likewise, we quantify a model’s Multilingual Ability (MA) based on its performance on (i) two question-answering tasks (i.e., MLQA Lewis et al. (2019) and XQuAD Artetxe et al. (2019)) and (ii) XLSUM Hasan et al. (2021), a summarization task.
1.2 Our Contributions
CFT Outperforms Mixture.
We first observe that models trained using our two-phase CFT setup perform better than the single-phase "dataset mixture" setup (Tables 12; §4.2). Moreover, our two-phase CFT setup overall results in a better model for all languages, including English, for the same number of training steps. The two-phase CFT pipeline also provides more flexibility than training on a mixture of datasets, with the possibility of extending our approach to multi-phase fine-tuning, especially when data from earlier phases might not be available.
Forgetting vs. Dataset Similarity.
As mentioned earlier, fine-tuning with multilingual datasets to enhance a model’s multilingual ability can lead to a decline in its English ability due to catastrophic forgetting (Mukhoti et al., 2023; Winata et al., 2023). We investigate the factors that may lead to such forgetting by computing the similarity of English and Multilingual Instruction Fine-tuning (IFT) datasets. We observe that when English and multilingual datasets have instructions that are not similar, there is a decline in the Phase 2 model’s performance in English. On the other hand, when Phase 1 and Phase 2 datasets encode similar instructions, the Phase 2 model’s performance in English improves (refer to Figure 1). To quantify the similarity of these phase-wise datasets, we introduce two metrics based on language-agnostic embeddings and model representations. We show that our quantification correlates with the decline in English ability (Tables 34; §4.3).
Mitigating Forgetting.
@@ 1.2 Our Contributions
We study the efficacy of two tailored variants of existing CFT strategies to mitigate the decline in EA after Phase 2 fine-tuning, while boosting MA. The first strategy is distribution replay. Here, we look at generative replay, i.e., using instructions from a similar English counterpart of the Phase 2 dataset to generate replay data using the Phase 1 model. We also try english replay which acts as language replay by utilizing existing English parallel data from the Phase 2 distribution. The second strategy employs layer freezing. Our heuristic selects specific layers for freezing during Phase 2 fine-tuning based on the weight differences between the Base and Phase 1 models. We also explore Spectrum Hartford et al. (2024) as an alternative heuristic. We study the gains in EA and MA of these strategies compared to specific baselines (Table 5; §5). To the best of our knowledge, we are the first to explore the effectiveness of CFT on LLMs with multilingual instruction datasets.
2 Related Work
Continual Learning in LLMs Chen et al. (2026).
In general, continual learning in LLMs can be broadly categorized into (i) continual pre-training (CPT) and (ii) continual fine-tuning (CFT). In CPT, the LLMs are continuously pre-trained to adapt to new domains or tasks by continuously updating them with new data alongside the existing data Shi et al. (2024). CPT builds on the existing LLM’s knowledge and is more computationally efficient than retraining an LLM using the current and old pre-training data Gupta et al. (2023). CPT is employed when distributional shifts occur (i) over time Amba Hombaiah et al. (2021); Jang et al. (2022a, b), (ii) across languages Jin et al. (2022); Fujii et al. (2024); Blevins et al. (2024) or (iii) across domains Ke et al. (2023); Gong et al. (2022); Xie et al. (2023).
On the other hand, CFT involves training the LLM on successive downstream tasks with varying data distribution or time shifts Shi et al. (2024). CFT comprises fine-tuning for different tasks Carrión and Casacuberta (2022); Guan et al. (2025), instruction-tuning Cahyawijaya et al. (2023); Kang et al. (2025), model refinement/editing Zhang et al. (2023) and alignment Suhr and Artzi (2023). Recent literature also focuses on using CFT to assist the LLM to learn new languages Praharaj and Matveeva (2023); Pfeiffer et al. (2022); Badola et al. (2023); Singh et al. (2024c).
CFT: Enhancing LLMs Multilingual Abilities.
Cahyawijaya et al. Cahyawijaya et al. (2023) propose InstructAlign which uses cross-lingual alignment and episodic replay to align an LLM’s pre-trained languages to unseen languages but requires parallel data and previous task data. Shaham et al. Shaham et al. (2024) introduces multilinguality during the first instruction fine-tuning phase which improves an LLM’s instruction following capability across languages. He et al. He et al. (2023) show catastrophic forgetting during CFT and use techniques such as joint fine-tuning and model regularization to mitigate it. However, these techniques are computationally expensive or require access to previous task data.
Multilingual Adaptation.
This set of works looks at language and task adaption by adjusting the model to understand new languages and enhancing its performance on specific tasks through fine-tuning, respectively Chen et al. (2023); Zhao et al. (2024); Pfeiffer et al. (2020). For instance, Chen et al. (2023) perform task adaption by fine-tuning the model on downstream task data. For language adaption, they fine-tune only the token embedding layer, helping the model learn specific lexical meanings of new languages. Language and english ability are either trained in parallel or sequentially. However, in this paper, we try to incorporate multilingual ability in models with the constraint that they may have already learned english ability (e.g., Mistral-7B-Instruct). To the best of our knowledge, this is a first attempt at studying the effect of task and language self-instruct datasets on an LLM’s multilingual ability through CFT.
3 Two-phase Continual Fine-tuning Setup
When instruction fine-tuning LLMs, the most natural method is to fine-tune on a "dataset mixture" containing English and Multilingual data Workshop et al. (2023). However, fine-tuning on all languages simultaneously may introduce performance bias where the model performs better in English (and other high resource languages) Dhamecha et al. (2021); Li et al. (2024a, b) [^2] In §4.2, we compare dataset mixture to CFT..
Continual Fine-tuning (CFT).
To improve the multilingual performance of pre-trained LLMs, we introduce the following two-phase CFT process.
@@ 3 Two-phase Continual Fine-tuning Setup
Two-Phase CFT Process Phase 1: Fine-tune a base LLM end-to-end on an English instruction dataset. Phase 1 aims to teach the LLM English Instruction Following Ability, which we refer to as English Ability (EA). Phase 2: Take the fine-tuned LLM from Phase 1 and further fine-tune it end-to-end on a Multilingual instruction dataset. Phase 2 focuses on enhancing the LLM’s Multilingual Ability (MA), using a dataset with multiple languages and fewer data points per language.
Challenges.
@@
The primary challenge in our two-phase CFT process is that the LLM’s Multilingual Ability must not come at the cost of its English Ability. We impose two additional constraints based on real-world scenarios. First, in Phase 2, we cannot re-use Phase 1’s dataset. Often instruction fine-tuned LLMs are available without their corresponding datasets (e.g., Mistral-7B-Instruct Jiang et al. (2023)). Second, in Phase 2, we cannot use the weights of the Phase 1 model during training, as saving both old and new set of parameters on the GPU for training would be computationally expensive.
@@
Two-phase Continual Fine-tuning Model Phase 1 (P1) Phase 2 (P2) IFEval $(\uparrow)$ Alpaca Eval $(\uparrow)$ MMLU $(\uparrow)$ HellaSwag $(\uparrow)$ XLSUM_en $(\uparrow)$ Average Dataset Dataset P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 P1 P2 Mistral-7B Alpaca 0.364 0.395 0.12 0.16 0.552 0.573 0.581 0.616 0.10 0.11 0.343 0.371 Instruct Multi 0.550 0.462 0.35 0.15 0.575 0.533 0.641 0.416 0.13 0.10 0.449 0.332 LLaMA-3-8B Alpaca Alpaca 0.277 0.326 0.10 0.11 0.231 0.242 0.556 0.567 0.07 0.08 0.247 0.265 Instruct 0.735 0.182 0.14 0.10 0.340 0.239 0.533 0.278 0.11 0.09 0.372 0.178 Dataset Mixture Model Dataset Mixture IFEval $(\uparrow)$ Alpaca Eval $(\uparrow)$ MMLU $(\uparrow)$ HellaSwag $(\uparrow)$ XLSUM_en $(\uparrow)$ Average Mistral-7B Alpaca MultiAlpaca 0.394 0.23 0.538 0.602 0.09 0.371 LLaMA-3-8B Alpaca MultiAlpaca 0.363 0.07 0.598 0.602 0.04 0.335
@@
Table 1: English Ability results for two-phase Continual Fine-tuning (CFT). When the phase-wise datasets are similar (Definition 1 and Definition 2), English Ability post Phase 2 (P2) fine-tuning consistently improves (denoted with green). When the phase-wise datasets are not similar, we see a significant decline in English Ability post Phase 2 (P2) fine-tuning (denote with red). We also provide numbers for dataset mixture – when the models are fine-tuned simultaneously on the Phase 1 and Phase 2 datasets.
4 Evaluating English & Multilingual Ability for Multilingual CFT
@@ 4.1 Experiment Setup & Evaluation Tasks
We continually fine-tune open-source Mistral-7B Jiang et al. (2023) and LLaMA-3-8B Dubey et al. (2024) LLMs for multilingual adaptation.
Fine-tuning Datasets.
@@
For our phase-wise datasets, we use the open-source Alpaca Taori et al. (2023), MultiAlpaca Wei et al. (2023), and OpenOrca Lian et al. (2023) datasets. Alpaca is a self-instruct English-only dataset. MultiAlpaca is a multilingual dataset created by translating Alpaca’s seed tasks to 11 languages and using GPT-3.5-Turbo for response collection. The languages are in equal proportions and are “French”, “Arabic”, “German”, “Spanish”, “Indonesian”, “Japanese”, “Korean”, “Portuguese”, “Russian”, “Thai”, and “Vietnamese”. The appendix (§A.2) describes OpenOrca and mOpenOrca.
Fine-tuning Technique.
We perform full fine-tuning in bf16 precision to study the impact of multilingual post-training in Phase 2 on English ability. This setting allows us to fully leverage model capacity, which may not be attainable under parameter-efficient fine-tuning methods Aggarwal et al. (2024); Panda et al. (2024). To mitigate the resulting degradation in English performance, §5 introduces a heuristic-based layer freezing strategy that selectively freezes a subset of layers while fine-tuning the remainder. All experiments are conducted using Axolotl333github.com/axolotl-ai-cloud/axolotl/, an open-source framework for large-scale LLM fine-tuning.
Evaluation Tasks.
@@
To quantify an LLM’s english ability, we evaluate Phase 1 and Phase 2 models on two instruction-following tasks (i) IFEval Zhou et al. (2023) and (ii) Alpaca Eval Li et al. (2023), (iii) MMLU Hendrycks et al. (2021) for problem-solving, (iv) HellaSwag Zellers et al. (2019) for commonsense reasoning ability, and (v) XLSUM_en Hasan et al. (2021) for summarization. To quantify an LLM’s multilingual ability, we evaluate our fine-tuned models on three benchmark datasets comprising two multilingual generative tasks: question answering (MLQA Lewis et al. (2019) & XQuAD Artetxe et al. (2019)), instruction-following (GMMLU Singh et al. (2024a)), and summarization (XLSUM Hasan et al. (2021)).
Our evaluation suite spans diverse linguistic phenomena and reasoning skills to comprehensively assess language understanding, generation, and problem-solving. It also incorporates parallel benchmarks for English and multilingual ability, enabling direct cross-lingual comparison. Further details are provided in §A.3.
@@
To evaluate our models on EA and MA, we use LM-Evaluation-Harness444github.com/EleutherAI/lm-evaluation-harness, which is a unified framework for zero/few-shot evaluations of LLMs. For both English and multilingual ability, we use zero-shot evaluation. For additional details on the training setup, code, and evaluation tasks, refer to §A.
4.2 Results